HUBBLE SPACE TELESCOPE DAILY REPORT #898 PERIOD COVERED: 0800 EST Daily Status Report as of 152/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 4725 (Scale Cycle 2 Calibration) The Fine Guidance Sensors were used to make an astrometric observations which will be used to calibrate the central region of the astrometric FGS. The observations was made during LOS, and remains of the Science Tape recorder. The data will be analyzed later. 1.2 Completed Three Sets of PC 3603 (Snapshot of Protoplanetary Nebulae) The Planetary Camera was used to make three images of the protoplanetary system W-Aquilae to study its current morphology. All three observations were received and displayed in the OSS. The first two images were reported as normal, and the last was reported as fine. 1.3 Completed Three Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) A grand total of 13 ultraviolet snapshot images were made with the Faint Object Camera (f/96) of three different nearby galaxies (ESO-1317-2, ESO-2333-3, and UGC-8256). All 13 images were reported to be satisfactory. 1.4 Completed PC 4770 (WF/PC Cycle 2 Calibration: Internal Monitor Proposal) The Planetary Camera was used to make one bias, four internal delta flat, and two dark calibration observations. The bias and internal flat observations were reported to be normal. The dark images were reported to have shown residual images from the K-spot observations. 1.5 Completed GHRS/2 4601 (Jupiter's Hydrogen Lyman Alpha Auroral/ Airglow Line Profiles: Cycle 3 Medium) The High Resolution Spectrograph as used to make "sky background" observations with will be used for calibration of observations of Hydrogen emission studies of Jupiter's atmosphere. A total of four observations were made, and they were reported to be satisfactory. 1.6 Completed FOC/96 4800 (The Stellar Content of Wolf - Rayet Galaxies (FOC Images)) The Faint Object Camera (f/96) was used to make ultraviolet images of the Wolf- Rayet galaxy POX4. This is a star bursting galaxy which is showing broad emission lines in Helium II at 4686 A0 from stellar sources. The observations were executed as scheduled, and no problems were reported. 1.7 Completed WFC 4381 (HST Medium - Deep Survey Cycle 3) The Wide Field Camera was used to make high latitude medium-deep sky survey images which were taken in parallel with the prime science observations. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.8 Completed FGS 2859 (Linking HIPPARCOS to the Extragalactic Reference Frame Part 2 of 6) Astrometric observations were made with Fine Guidance Sensor #3 on the extragalactic object 1758-651IN and on a field of nearby stars. The observations were executed as scheduled, and monitored in the OSS. No problems were reported. 1.9 Completed FGS 4031 (High Speed Astrometry - A search for Planetary Companions to Low Mass) The Fine Guidance Sensors were used to make astrometric observations of the star Proxima Centauri to look for a Jupiter sized planet in orbit around the target star as measured from the position perturbations caused by the planet. The observations were executed as scheduled, and no problems were reported. 1.10 Completed FOC/96 4152 (FOC Observations to Monitor Distortion) The Faint Object Camera (f/96) was used with filter F486N to monitor the focus of the HST OTA looking for changes due to desorption. A total fo five images were taken. A corrective offset maneuver was calculated and sent to the spacecraft after the first image. All subsequent images were reported to be normal. 1.11 Completed FOC/96 4153 (FOC UV Throughput Monitoring) The Faint Object Camera was used to make five observations of the standard star PMB16274 to determine the ultraviolet throughput of the HST-FOC. The first image was received and displayed. Based on its contents a corrective maneuver was calculated and uplinked to the spacecraft. The remaining observations were reported to be satisfactory. 1.12 Completed WFC/PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) Both the Wide Field Camera and the Planetary Camera were used to make internal flat images looking for new forms of contamination (i.e., daises and Measles), and the possible return of quantum efficiency hysteresis. Both observations were received and displayed in the OSS. Both were reported to be satisfactory. 1.13 Completed Two Sets of PC 4929 (WF/PC cycle 2: Improved Dark Calibration) The Planetary Camera was used to make two sets of two each internal dark calibration images. All four images were 1200 second long exposures, and all were preflashed. All four observations were reported to be satisfactory. 1.14 Completed FOS/BL 3683 (Accretion Disk Mapping in Eclipsing Cataclysmic variables) The Faint Object Spectrograph was used to make rapid mode observations of the cataclysmic variable star IP Pegasi during the eclipse with its accretion disk. The onboard target acquisition was successful. The spectral observations were monitored and appear to be very successful. 1.15 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.16 Completed GHRS/96 4569 (A Search for Microflaring Activity in Dwarf M Stars - Cycle 3 High) The High Speed Photometer was used to make observations of the dwarf star CN Leonis looking for microflares which have short duration and occur with numbers inversely proportional to the flares energy. The onboard target acquisitions left the target well centered. The scientific observations were received in the OSS. No anomalies were noted. 1.17 Completed GHRS/2 3444 (Composition of Gas in Individual Interstellar Clouds: Cycle 2 Augmentation) The High Resolution Spectrograph was used to make spectral observations of early type star HD215733. The target acquisitions were executed without problems. The wavelenght calibration and the science observations were received and displayed. They all were reported to look okay. 1.18 Completed WFC 4277 (Calibration of Nearby Galaxy Type IA Supernova as Standard Candles) The Wide Field Camera was used to make two images of galaxy NGC-5253 to help determine the periods of Cepheid variable stars contained within the galaxy. This will make it possible to accurately determine the distance to this galaxy. Two brief recentering events were reported to have occurred during the first image. No anomalies were reported during the second image. 1.19 Completed FOS/BL 4222 (Mass Exchange Binaries (FOC 34) - Cycle 1) The Faint Object Spectrograph (Blue Side) was used to make spectral observations of the binary star HZ Herulis. This is a binary system having a normal star with a contact companion. The observations were made in the rapid mode using grating G160L to learn more about the soft X-ray flux from the compact star. The OSS monitored the observations and assessed the data quality. No problems were reported. 1.20 Completed Two Sets of PC 3848 (Separate Light ?curves for Pluto and Charon) The Planetary Camera was used to make a total of four images of the planet Pluto and its moon Charon individually. All four images were reported to look fine. 1.21 Completed FOC/96 4846 (High Resolution Observations of Cataclysmic Variables: Cycle 3) The Faint Object Camera (f/96) was used to make two images of the complex regions of nebulosity surrounding Nova Cygni 1992. A couple of recentering events occurred during the observations, however, no anomalies were reported in the images. The observations were displayed for the principal investigator, and several hard copies were made. 1.22 Completed WFC 4848 (GTO Low Latitude Parallel Program Cycle 3 High) The Wide Field Camera was used to make a series of low latitude observations taken in parallel with the prime science observations. There was a recentering event reported during the observations, but no problems were noted with the telemetry. The observations were received and displayed, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 26 Successful : 26 Scheduled Reacquisitions : 22 Successful : 22 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 53 Successful : 51 149/2306Z, HSTAR-4103, a roll delay update to correct for accumulated V1 errors. Subsequent acquisitions A-Okay. 151/1758Z, HSTAR-4107 Update trackers 2 & 3 failed, Tracker 2 failed all three iterations. subsequent Acquisitions A-Okay. 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOC/96 3801 (Search for QSOs Suitable for Subsequent Observations of He II) 3.2 Complete GHRS/2 4601 (Jupiter's Hydrogen Lyman Alpha Auroral/Airglow Line Profile) 3.3 Complete Three Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) 3.4 Complete PC 3887 (Integrated Dynamical and Spectroscopic Observations of Jupiter and Satellites) 3.5 Complete WFC/PC 4109 (WF/PC cycle 2 Calibrations: Rapid Internal Monitor) 3.6 Complete PC 3603 (Snapshots of Protoplanetary Nebulae) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-